ng Kentucky University-CHEM 120-Spring18-STEVENS Activities and Due Dates HW 10 Resources 5/9/20181 159 PM O 70.7/100?5/7/2018 0834 PM O Assignm Queston 10 of 25 Available F Due Date: Late submi The vapor pressan, Pdacertain kud was measured at two temperatures If you were going to graphically determine the enthalpy of vaporizaton,f tis liquid, what points would you plot? The data is shown in this table. Number TK) P(kPa) 225 396 To avoid rounding errors, use three point 1 Points Possi Grade Cateç Description Policies: significant figures in x values and four signaficant figures in he y values NOTE: Keep the pressure units in kPa point 2:10 You can che You can view give up an a You have ter You lose 10% answer in ya attempt at th For multiple depends on t Determine the rise, run, and slope of the line formed by these points. rise siope Number What is the enthalpy of vaporization of this liquid? Number AFT J/ mol DeTextbook O Previous Give Up& View Solution Cheak Answer NeatExtHelp With TI ?? r10 112
Answer #1

here graph is drawn between ln K   and 1/T

lnK = y -axis

1/T = x - axis

point 1:   x = 4.44 x 10^-3   ,   y = 1.376

point 2 : x = 1.60 x 10^-3   , y = 1.864

rise =   y2-y1 = 0.488

run = x2 - x1    = - 2.84 x 10^-3

slope = rise / run = - 0.488 / 2.84 x 10^-3 = - 171.8

-H / R = slope

H = 171.8 x 8.314 = 1429 J / mol